Solve the following equation for y (get y by itself on one side of the equal sign): 5x - 3y = 15
Found 2 solutions by LarissaRichardson, Alan3354:
Answer by LarissaRichardson(34)

5x - 3y = 15
5x-15=3y
Divide by 3
5x-15/3=y

Answer by Alan3354(69443)

5x - 3y = 15
-3y = -5x + 15
y = (5/3)x - 5